# Quantum-limited optical delay sensing across an enhanced dynamic range by frequency-resolving two-photon interference

### v1

@@ Abstract
Hong–Ou–Mandel interferometry is an effective tool for precision sensing, but it faces a fundamental trade-off between dynamic range and measurement precision. To overcome this limitation, we sample the frequencies of two interfering photons, which extends the sensitivity in the time domain. This enhances the dynamic range of optical delay estimation by up to twenty times compared to the non-resolved estimates. We demonstrate this approach with independent photon sources and show the precision scaling with the detector frequency resolution. This technique enables scan-free nanometre resolution depth sensing over a millimetre-scale range, with applications in biological and nanomaterial imaging.
Two-photon interference is a fundamental effect in quantum optics that underpins applications in optical quantum computing, communication, and sensing [1, 2, 3, 4, 5, 6, 7, 8]. In sensing, interference between single photons can resolve microscopic path length differences with high sensitivity. The Hong-Ou-Mandel experiment first established this technique for time delay estimation and demonstrated sub-picosecond timing precision [9]. Combining estimation theory and source engineering has enabled the precision to reach attosecond scale [10, 11, 12]. In comparison, direct time-of-flight using single-photon detectors is limited by detector jitter on the order of 10s picoseconds [13].
Unlike classical interferometry, two-photon interference arises from the coherent sum of two-photon probability amplitudes [9, 5]. It requires indistinguishability in all degrees of freedom, including arrival time, rather than first-order coherence, which can be challenging stabilise [14, 15]. This makes two-photon interference robust to phase noise and imbalanced losses, which significantly impact the measurement precision of phase-sensitive interferometers [16].
A two-photon interferogram has a distinct single fringe centred on zero delay between the interfering photons. This fringe has a width approximated by the photon coherence time $\tau_{c}\sim 1/\Delta\omega$, where $\Delta\omega$ is the photon spectral bandwidth. It enables unambiguous delay estimation over a much wider range than phase-based interferometers that produce periodic fringes of width $\lambda$, the optical wavelength, and cause phase wrapping ambiguity beyond one cycle [17, 18]. In both cases, narrower fringes increase the sensitivity to optical path length but reduce the unambiguous range, revealing a fundamental trade-off between sensitivity and dynamic range [19, 20]. We aim to overcome this trade-off by resolving the photonic frequency modes.
@@
Frequency-resolved sampling can access more information between interfering photons [21] and restore interference between temporally distinguishable photons [22, 23, 24]. Recently, this technique was shown theoretically to achieve quantum-limited sensitivity for time delay estimation, saturating the quantum Cramér-Rao bound, provided photons remain perfectly indistinguishable [7, 25]. The frequency-resolved method offers nearly uniform sensitivity across the entire delay range, unlike the non-resolved method, which is parameter-dependent and must be optimised to reach maximum sensitivity [10]. The interference visibility depends on the spectral resolution, which must satisfy $\delta\omega\ll 1/\lvert\Delta t\rvert$ to restore interference at delay $\Delta t$ [7]. This has been studied using a spectrally detuned entangled source [26], and weak coherent states [27].
@@
In this work, we focus on the metrological advantage of the frequency-resolved estimates compared to the unresolved measurement using independent, spectrally uncorrelated single-photon states. We employ high-resolution ($6.9/2\pi$ GHz) frequency-resolved detectors, benchmark this method against a maximum-likelihood estimator without spectral information and realise more than a twenty-fold increase in dynamic range and a thousand-fold improvement relative to other implementations [10, 11, 12]. We analyse the effects of finite frequency resolution on the Fisher information and discuss the practical utility of this technique for precision depth sensing.
@@
where $\Delta\omega=\omega_{1}-\omega_{2}$ is the difference in frequency and $\mathcal{C}(\Delta\omega)$ is the joint spectral distribution, see SM0.9. For a fixed delay $\Delta t$, the output distributions in Eqs. 1 and 2 can be sampled by measuring both the bunching ($N_{B}$) and coincidence ($N_{C}$) outcomes. In addition, sampling the joint frequencies $(\omega_{1},\omega_{2})$ is used to find the frequency difference $\Delta\omega$. The time delay $\Delta t$ is then inferred through maximum likelihood estimation (MLE), see LABEL:SM_theory.
@@
In contrast, the frequency-resolved distribution given in Eq. 2 does not yield an analytic solution for $\Delta t$. Instead, we construct the estimator numerically via the likelihood function. Let the sample of $N$ measured outcomes be denoted by $S_{N}=\{s^{\{i\}}\}_{i=1}^{N}$, where each $s^{\{i\}}=(\Delta\omega^{\{i\}},\delta^{\{i\}})$ contains a frequency difference and bunching/anti-bunching outcomes. The likelihood of a delay value $\Delta t$ given the sample is defined as $\mathcal{L}_{R}(\Delta t\mid S_{N})=P(S_{N}\mid\Delta t)$. The estimator $\widetilde{\Delta t}_{\text{MLE}}$ is the value with the largest log-likelihood, see SM0.4,
@@
The only unknown parameter remaining in Eq. (2) is $\eta$, which can be obtained experimentally in the calibration step. Therefore, the frequency-resolved scheme provides a numerical estimate of the time delay by matching the measured outcomes to the theoretical distribution via maximum likelihood. The error on these parameters will feed forward into the variance of the estimator which we discuss in SM0.5. A full derivation of the estimator in Eq. (4) is provided in Ref. [7].
We perform the experiment using two configurations. The first uses heralded single photons from two independent parametric down-conversion (PDC) sources, each with Gaussian spectra and without spectral correlations. The second configuration uses a single PDC source with higher pair rates and indistinguishability, see SM0.1. Each photon-pair source consists of a domain-engineered, aperiodically poled KTP crystal [29, 30], designed for group-velocity matching with a pulsed pump laser. This enables the generation of spectrally pure, mutually indistinguishable photon pairs at $1550\text{\,}\mathrm{nm}$ [31, 32], see supplement for rates Quantum-limited optical delay sensing across an enhanced dynamic range by frequency-resolving two-photon interference.
@@
The precision of the experimental estimates against the theoretical bound is calculated from the experimental metrological information, $\mathscr{F}_{exp}$, using the variance of the estimates via the Cramér–Rao bound [33, 34] (CRB), $\mathrm{var}[\widetilde{\Delta t}]\geq 1/N\mathscr{F}_{exp}$, where $N$ is the number of samples per estimate. Fig. 4 presents both the experimentally obtained metrological information and the theoretical Fisher information across the full delay range for the non-resolved (a) and frequency-resolved (b) estimates. The Fisher information of the non-resolved estimate (solid line) is calculated using Eq. 6, with the maximum photon indistinguishability $\eta=0.98$ obtained from calibration, see SM0.1. The Fisher information peaks at 0.336 $\text{ps}^{2}$ and falls off with a width approximately $2\tau\approx 2$ ps. The experimental data align well with the theoretical prediction, confirming the loss of precision at delays greater than $2\tau$.
In frequency-resolved metrological information is plotted in Fig. 4(b) using Eq. 7, where both $\eta$ and $\Delta t$ vary, and the values of $\eta$ are found from Fig. 2(c). The experimental values agree closely with theory for the 2-fold configuration. However, the 4-fold estimates achieve a metrological lower Fisher information than expected. This reduced performance may result from the increased complexity of the heralded setup, including additional noise from fibre paths and the cumulative effect of timing jitter across twice as many detectors.
@@
Figure 4: The experimental metrological information $\mathscr{F}_{exp}$ is plotted against the estimated delay as solid points for 2-fold and empty points for 4-fold estimates. (a) The non-resolved Fisher information is plotted as a solid orange line using Eq. (6) and $\eta=0.98$, displaying a peak at around 1 coherence length, and the experimental points agree well. The inset shows the ratio of the non-resolved and resolved mean squared error (MSE) over the same delay range in picoseconds. (b) The finite frequency-resolved Fisher information is plotted as a solid blue line using Eq.7 and shows the extended sensitivity range. The red crosses are estimates using a lower pump power (100 mW) and improved indistinguishability. The grey horizontal line is the QFI with value 0.52 ($\text{ps}^{-2}$).
Resolving the frequency difference between interfering photons greatly enhances the dynamic range of optical delay estimation relative to the identical non-resolved method. The achievable precision depends on both the spectral bandwidth and the resolution of the frequency-resolving detectors, which can be improved in future experiments. We derived the Fisher information for finite frequency resolution in SM0.10. Improving the resolution by a factor of ten extends the dynamic range also by a factor of ten, and an increase by a factor of sixty nearly saturates the quantum Cramér-Rao bound over the full 60 ps delay range. Compared with a recent experiment that reached attosecond precision  [10], we obtained a similar per-sample performance even after incorporating their loss model. The best measured precision in Ref. [10] was 4.7 attoseconds using 414 billion samples, while we reached 500 attoseconds using 7.9 million samples. The usual trade-off between precision and dynamic range appears to be avoided for delays up to 10 ps with the frequency-resolved method. In this regime, the Fisher information exceeds the peak of the non-resolved Fisher information, so the need for careful optimisation near that peak is removed while retaining equal or greater precision.
The main limitation of our present setup is loss from chromatic dispersion using standard single-mode fibres for time-of-flight spectrometry, which is about 99% for 50 km spools. This can be reduced by using commercial single-photon spectrometers with lower loss, a chirped fibre Bragg grating [22], or a diffraction grating combined with a SPAD array [25, 27]. However, the resolution from these devices is typically below that of time-of-flight spectroscopy with dispersive fibres. The precision of our setup can be further enhanced using entangled photon pairs to realise a $\sqrt{2}$ improvement per sample [25]. Further gains are possible with spectrally engineered sources that use detuned bi-photon schemes [12, 11], which can provide large Fisher information without elaborate fringe tracking or calibration.
@@ Acknowledgements
SM0.1 Experimental Setup
A detailed experimental layout of the independent source configuration is shown in Fig. SM1. The experiment employs two configurations: (I) heralded photons from independent sources, and (II) a single photon-pair source. In configuration (I), two independent photon-pair sources (S1 and S2) are pumped by a Ti:sapphire laser. Each source generates a photon pair, with one photon from each pair separated using a polarising beam splitter (PBS) and coupled into a single-mode fibre (SMF), where it serves as a herald. Long-pass silicon windows are used to filter out the pump light. The remaining photons are also coupled into SMFs and launched onto a balanced 50:50 beam splitter (BS). Each photon input ot the BS is prepared with horizontal polarisation using a linear polariser and in-fibre polarisation controllers. The output ports of the BS are each connected to a 50 km SMF spool, forming a time-of-flight spectrometer. These outputs are then split again into two fibres, forming pseudo-number-resolving detectors. Photons are detected using superconducting nanowire single-photon detectors (SNSPDs) with approximately 80% quantum efficiency and 50 ps timing jitter.
@@
SM0.2 Frequency-resolving detector
The time-of-flight (ToF) spectrometer is made up of 50 km fibre spool with a total group delay dispersion (GDD) $\phi^{\prime\prime}=$909\text{\,}\mathrm{p}\mathrm{s}^{2}$\$, combined with detector jitter, resulting in a spectral resolution $\delta\omega$ of $6.9/2\pi$ GHz. The frequency is given by $(\omega^{i}-\Omega)=(t^{i}-t_{\Omega})/\phi^{\prime\prime}$, where $t^{i}$ is the arrival time of the $i^{th}$ photon triggered by a sync signal from our Ti:Sapph laser and $\Omega=$193\text{\,}\mathrm{THz}$\$ is the central frequency of our 1550 nm photons. Since only the frequency difference $\Delta\omega^{i}$ is used in the MLE, we can set $t_{\Omega}=0$. The frequency of both output ports is then resolved as a linear function of the “time-of-flight” $\Delta\omega^{i}=\Delta t^{i}/\phi^{\prime\prime}$ through fibre with a dispersion factor $1/\phi^{\prime\prime}=0.0011(4)$ nm/ps.
The joint spectral intensity figures Using the frequency-resolved detectors
SM0.3 Calibration and experimental parameters
@@
SM0.4 Estimation Protocol
@@
SM0.5 Estimator noise
Let the measured frequency shifts be $\{\Delta\omega_{i}\}_{i=1}^{N}$ with true density $P_{n}(\Delta\omega)$. We estimate $\Delta t$ by maximising the log likelihood $\ell(\Delta t)=\sum_{i=1}^{N}\log P(\Delta\omega_{i}\mid\Delta t)$, which is equivalent to using $P(\Delta t\mid\Delta\omega_{i})$ up to a normalising constant.
Assume a small model mismatch between the true density $P_{n}$ and the model $P(\,\cdot\,\mid\Delta t)$. Write $\Delta\omega_{i}=\tilde{\Delta\omega}_{i}+\varepsilon_{i}$, where $\tilde{\Delta\omega}_{i}\sim P(\,\cdot\,\mid\Delta t)$ and $\varepsilon_{i}$ captures the difference between the two distributions with $\mathbb{E}[\varepsilon_{i}]=0$ and $\mathbb{E}[\varepsilon_{i}^{2}]=\sigma_{\varepsilon}^{2}$.
A second-order Taylor expansion of each log likelihood term about $\tilde{\Delta\omega}_{i}$ gives
$\log P(\Delta\omega_{i}\mid\Delta t)=\log P(\tilde{\Delta\omega}_{i}\mid\Delta t)+\partial_{\Delta\omega}\log P(\tilde{\Delta\omega}_{i}\mid\Delta t)\,\varepsilon_{i}+\tfrac{1}{2}\,\partial_{\Delta\omega}^{2}\log P(\tilde{\Delta\omega}_{i}\mid\Delta t)\,\varepsilon_{i}^{2}+o(\varepsilon_{i}^{2}).$
Taking expectations over the mismatch, the first order term vanishes since $\mathbb{E}[\varepsilon_{i}]=0$. The second order term contributes a bias proportional to $\mathbb{E}[\varepsilon_{i}^{2}]$, hence to $\sigma_{\varepsilon}^{2}$. This contribution increases as $P_{n}$ and $P(\,\cdot\,\mid\Delta t)$ diverge.
In the regime $|\Delta\omega\,\Delta t|\ll 1$, the two distributions are practically equivalent, so $\sigma_{\varepsilon}^{2}$ is negligible and the induced bias can be ignored to an excellent approximation.
SM0.6 Spectral indistinguishability between one and two sources
@@
SM0.7 Stability of the setup over time
Collecting many samples over a long period of time can cause bias in the estimates due to experimental drift. One way to mitigate this is to analyse the Allan variance of the experimentally obtained samples for the bunching and coincidence outcomes, $\alpha(X)$, where X=1,1X=1,-1 and the spectral differences. In Fig. SM3 we have plotted the Allan variance for the BoC and spectral differences, both the calculated using the following form,
$\sigma_{y}^{2}(\tau)=\frac{1}{2\,(M-1)}\sum_{j=1}^{M-1}\bigl(\tilde{y}_{j+1}-\tilde{y}_{j}\bigr)^{2}.$ (SM2)
@@
SM0.8 Maximum resolution estimation
In this section, we analyse the resolution of the resolved optical delay estimation using our best prepared state. We use low pump power (100 mW) and achieve HOM visibility of 0.98, which improves the estimate precision. Fig. SM4 displays the cumulative estimate for the delays prepared at 6.567 ps and 6.570 ps.
SM0.9 Derivation of Probability function in terms of frequency difference
@@
$P_{\eta}(\omega,\omega^{\prime},X)=\gamma^{2}\xi(\omega)^{2}\xi(\omega^{\prime})^{2}\frac{1+\alpha(X)\eta^{2}\cos\left((\omega-\omega^{\prime})\Delta t\right)}{2},$ (SM3)
@@
$\xi(\omega)^{2}=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left[-\frac{\left(\omega-\Omega_{0}\right)^{2}}{2\sigma^{2}}\right].$ (SM4)
@@
$P_{\eta}(W,\Delta\omega,X)=\gamma^{2}\xi^{2}\left(W+\frac{\Delta\omega}{2}\right)\xi^{2}\left(W-\frac{\Delta\omega}{2}\right)\frac{1+\alpha(X)\eta^{2}\cos\left(\Delta\omega\Delta t\right)}{2}.$ (SM5)
@@
$\begin{split}\xi^{2}\left(W+\frac{\Delta\omega}{2}\right)\xi^{2}\left(W-\frac{\Delta\omega}{2}\right)&=\frac{1}{2\pi\sigma^{2}}\exp\left[-\frac{(W-\Omega_{0}+\Delta\omega/2)^{2}}{2\sigma^{2}}\right]\exp\left[-\frac{(W-\Omega_{0}-\Delta\omega/2)^{2}}{2\sigma^{2}}\right]\\ &=\frac{1}{2\pi\sigma^{2}}\exp\left[-\frac{(W-\Omega_{0})^{2}}{\sigma^{2}}\right]\exp\left[-\frac{\Delta\omega^{2}}{4\sigma^{2}}\right],\end{split}$ (SM6)

### v2

@@ Abstract
Optical sensing schemes that rely on two-photon interference provide a powerful platform for precision metrology, although they are inherently constrained by a trade-off between dynamic range and measurement precision. To overcome this limitation, we sample the frequencies of two interfering photons, which extends the sensitivity in the time domain. This enhances the dynamic range of optical delay estimation by up to twenty times compared to the non-resolved estimates. We demonstrate this approach with independent photon sources and show the behaviour of finite frequency resolving detectors. This technique enables scan-free nanometre resolution depth sensing over a millimetre-scale range, with applications in biological and nanomaterial imaging.
Two-photon interference is a consequence of the quantum nature of the electromagnetic field, and is foundational to many quantum optics experiments, from photonic quantum computing and communication to metrology and sensing [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Indistinguishable photons impinging on a balanced beam splitter manifest the non-classical tendency to exit through the same output port from the destructive interference of the two-photon probability amplitudes [1, 2, 13, 3]. This phenomenon has been exploited for precision estimation of optical path delays and coherence measurements, typically achieving sub-picosecond [14, 15, 16, 17] and recently attosecond temporal precision [18, 19, 20]. In comparison, direct time-of-flight using single-photon detectors is limited by detector jitter on the order of tens of picoseconds [21, 22]. However, methods that rely on measuring only the photon statistics at the output of a beam splitter are limited by the photons’ temporal bandwidth, which constrains the operational range and prevents sensitive measurements once the delay exceeds the coherence time [18, 23, 24].
Conventional multiphoton interference experiments do not resolve the photonic inner modes, such as frequency and time, which ignores the useful quantum information encoded within the photons [9, 25]. It was shown theoretically in Ref. [25] that resolving the photon frequency can restore interference even between non-identical photons and was later experimentally demonstrated between temporally distinguishable photons [26]. Similarly, frequency-resolved two-photon interference reveals quantum beat notes in the interferogram [27, 28, 29], in contrast to the single fringe of the non-resolved interferogram [1], accessing the spectral correlations that can be harnessed for precision metrology.
@@
Recently, metrological schemes that sample the photonic inner modes have been shown to achieve the ultimate quantum-limited sensitivity in estimating conjugate variables, including time–frequency shifts [8, 30], transverse position–momentum displacements [31, 32], and full-state polarisation [33]. In optical delay estimation, ideal frequency-resolved sampling offers uniform sensitivity across the entire delay range, unlike the non-resolved method, which is parameter-dependent and must be optimised to reach maximum sensitivity [18]. However, these theoretical works require infinite resolving capabilities to fully erase information in the conjugate space and restore quantum interference [34, 35]. In the absence of such devices, experiments are limited by finite resolution. The interference visibility depends on the spectral resolution, which must satisfy $\delta\omega\ll 1/\lvert\Delta t\rvert$ to restore interference with temporal delay $\Delta t$ [8]. This has been studied using non-degenerate spectrally entangled photons [36], and weak coherent states [37].
@@
In this work, we report on the metrological advantage of frequency-resolved sampling for optical delay estimation with independent, spectrally uncorrelated single-photon states. We employ high-resolution ($6.9/2\pi$ GHz) frequency-resolved detectors, benchmark this method against a maximum-likelihood estimator without the spectral information and realise more than a twenty-fold increase in dynamic range and a thousand-fold improvement relative to other implementations [18, 19, 20]. We analyse the effects of finite frequency resolution on the Fisher information and discuss the practical utility of this technique for precision depth sensing.
@@
where $\Delta\omega=\omega_{1}-\omega_{2}$ is the difference in frequency and $\mathcal{C}(\Delta\omega)$ is the joint spectral distribution, see SM9. For a fixed delay $\Delta t$, the output distributions in Eqs. 1 and 2 can be sampled by measuring both the bunching ($N_{B}$) and coincidence ($N_{C}$) outcomes. In addition, sampling the joint frequencies $(\omega_{1},\omega_{2})$ is used to find the frequency difference $\Delta\omega$. The time delay $\Delta t$ is then inferred through maximum likelihood estimation (MLE), see SM9.
@@
In contrast, the frequency-resolved distribution given in Eq. 2 does not yield an analytic solution for $\Delta t$. Instead, we construct the estimator numerically via the likelihood function. Let the sample of $N$ measured outcomes be denoted by $S_{N}=\{s^{\{i\}}\}_{i=1}^{N}$, where each $s^{\{i\}}=(\Delta\omega^{\{i\}},\delta^{\{i\}})$ contains a frequency difference and bunching/anti-bunching outcomes. The likelihood of a delay value $\Delta t$ given the sample is defined as $\mathcal{L}_{R}(\Delta t\mid S_{N})=P(S_{N}\mid\Delta t)$. The estimator $\widetilde{\Delta t}_{\text{MLE}}$ is the value with the largest log-likelihood, see SM4,
@@
The only unknown parameter remaining in Eq. (2) is $\eta$, which can be obtained experimentally in the calibration step. Therefore, the frequency-resolved scheme provides a numerical estimate of the time delay by matching the measured outcomes to the theoretical distribution via maximum likelihood. The error on these parameters will feed forward into the variance of the estimator which we discuss in SM5. A full derivation of the estimator in Eq. (4) is provided in Ref. [8].
We perform the experiment using two configurations. The first uses heralded single photons from two independent parametric down-conversion (PDC) sources, each with Gaussian spectra and without spectral correlations. The second configuration uses a single PDC source with higher pair rates and indistinguishability, see SM1. Each photon-pair source consists of a domain-engineered, aperiodically poled KTP crystal [38, 39], designed for group-velocity matching with a pulsed pump laser. This enables the generation of spectrally pure, mutually indistinguishable photon pairs at $1550\text{\,}\mathrm{nm}$ [40, 41], see supplement for rates Quantum-limited optical delay sensing across an enhanced dynamic range by frequency-resolving two-photon interference.
@@
The precision of the experimental estimates against the theoretical bound is calculated from the experimental metrological information, $\mathscr{F}_{exp}$, using the variance of the estimates via the Cramér–Rao bound [42, 43] (CRB), $\mathrm{var}[\widetilde{\Delta t}]\geq 1/N\mathscr{F}_{exp}$, where $N$ is the number of samples per estimate. Fig. 4 presents both the experimentally obtained metrological information and the theoretical Fisher information across the full delay range for the non-resolved (a) and frequency-resolved (b) estimates. The Fisher information of the non-resolved estimate (solid line) is calculated using Eq. 6, with the maximum photon indistinguishability $\eta=0.98$ obtained from calibration, see SM1. The Fisher information peaks at 0.336 $\text{ps}^{2}$ and falls off with a width approximately $2\tau\approx 2$ ps. The experimental data align well with the theoretical prediction, confirming the loss of precision at delays greater than $2\tau$.
In frequency-resolved metrological information is plotted in Fig. 4(a) using Eq. 7, where both $\eta$ and $\Delta t$ vary, and the values of $\eta$ are found from Fig. 2(c). The experimental values agree closely with theory for the 2-fold configuration. However, the 4-fold estimates achieve a metrological lower Fisher information than expected. This reduced performance may result from the increased complexity of the heralded setup, including additional noise from fibre paths and the cumulative effect of timing jitter across twice as many detectors.
Figure 4: (a) The experimental metrological information $\mathscr{F}_{exp}$ for the 2-fold estimates with frequency-resolved (FR) and non-resolved (NR) measurements is plotted against the prepared delay. The non-resolved Fisher information is plotted as a solid orange line using Eq. (6) and $\eta=0.98$, displaying a peak at around 1 coherence length, while the frequency-resolved Fisher information is plotted as a solid blue line using Eq.7 and displays an extended sensitivity range. The blue dashed horizontal line is the QFI with value 0.52 ($\text{ps}^{-2}$). (b) The metrological advantage is plotted for each estimate in logarithmic scale using the ratio of the resolved and non-resolved mean squared error, $\mathrm{MSE_{NR}/MSE_{FR}}$.
@@
Resolving the frequency difference between interfering photons greatly enhances the dynamic range of optical delay estimation relative to the identical non-resolved method. The achievable precision depends on both the spectral bandwidth and the resolution of the frequency-resolving detectors, which can be improved in future experiments. We derived the Fisher information for finite frequency resolution in SM10. Improving the resolution by a factor of ten extends the dynamic range also by a factor of ten, and an increase by a factor of sixty nearly saturates the quantum Cramér-Rao bound over the full 60 ps delay range. Compared with a recent experiment that reached attosecond precision  [18], we obtained a similar per-sample performance even after incorporating their loss model. The best measured precision in Ref. [18] was 4.7 attoseconds using 414 billion samples, while we reached 500 attoseconds using 7.9 million samples. The usual trade-off between precision and dynamic range appears to be avoided for delays up to 10 ps with the frequency-resolved method. In this regime, the Fisher information exceeds the peak of the non-resolved Fisher information, so the need for careful optimisation near that peak is removed while retaining equal or greater precision.
The main limitation of our present setup is loss from chromatic dispersion using standard single-mode fibres for time-of-flight spectrometry, which is about 99% for 50 km spools. This can be reduced by using commercial single-photon spectrometers with lower loss, a chirped fibre Bragg grating [26], or a diffraction grating combined with a SPAD array [44, 37]. However, the resolution from these devices is typically below that of time-of-flight spectroscopy with dispersive fibres. The precision of our setup can be further enhanced using entangled photon pairs to realise a $\sqrt{2}$ improvement per sample [44]. Further gains are possible with spectrally engineered sources that use detuned bi-photon schemes [20, 19], which can provide large Fisher information without fringe tracking or calibration, overcoming the sensitivity trade-off in optical sensors [45, 46].
Another advantage of two-photon interference is the robustness to phase noise and imbalanced losses, which significantly impact the measurement precision of phase-sensitive interferometers [23]. The sensitivity of our approach is also comparable to phase-based methods but without active stabilisation [2, 13] or the phase wrapping problem, which requires additional phase tracking algorithms [47, 48].
@@ Acknowledgements
SM1 Experimental Setup
A detailed experimental layout of the independent source configuration is shown in Fig. SM1. The experiment employs two configurations: (I) heralded photons from independent sources, and (II) a single photon-pair source. In configuration (I), two independent photon-pair sources (S1 and S2) are pumped by a Ti:sapphire laser. Each source generates a photon pair, with one photon from each pair separated using a polarising beam splitter (PBS) and coupled into a single-mode fibre (SMF), where it serves as a herald. Long-pass silicon windows are used to filter out the pump light. The remaining photons are also coupled into SMFs and launched onto a balanced 50:50 beam splitter (BS). Each photon input on the BS is prepared with horizontal polarisation using a linear polariser and in-fibre polarisation controllers. The output ports of the BS are each connected to a 50 km SMF spool, forming a time-of-flight spectrometer. These outputs are then split again into two fibres, forming pseudo-number-resolving detectors. Photons are detected using superconducting nanowire single-photon detectors (SNSPDs) with approximately 80% quantum efficiency and 50 ps timing jitter.
@@
Figure SM1: Experimental layout with two sources. A 775 nm pulsed Ti:Sapphire laser with 1.3 ps duration and 80 MHz repetition rate pumps two photon pair sources. The sources comprise of a aperiodicity-poled KTP crystal housed in a Sagnac configuration, generating spectrally degenerate Gaussian profile photon pairs at 1550 nm. One photon from each pair is separated using dichroic polarising beam splitters (dPBS) and coupled to a single-mode fibre (SMF) to be detected imediately on superconducting nanowire detectors (SNSPD) and time-stamped on Hydra harp time tagger. The other pair from each source is then coupled to SMF and launched into free space onto a 50:50 beam splitter. The polarisation of each independent photon is prepared using a linear plate polariser set to Horizontal to match the polarisation degree of freedom. The Spectral component is also matched by calibrating the temperature of the crystal ovens to yield the highest independent-HOM visibility, found to be 91%. Each path exiting the beam splitter is coupled to 50km SMF to induce chromatic dispersion and delay the photon packets in time. A time-of-flight spectrometer converts the time difference between the arrival of a photon and the arrival of the Ti:Sapphire sync pulse to estimate the frequency component after experience dispersion. Finally each path is multiplexed into two spatial modes to act as a pseudo number-resolving detectors.
SM2 Frequency-resolving detector
The time-of-flight (ToF) spectrometer is made up of 50 km fibre spool with a total group delay dispersion (GDD) $\phi^{\prime\prime}=$909\text{\,}\mathrm{p}\mathrm{s}^{2}$\$, combined with detector jitter, resulting in a spectral resolution $\delta\omega$ of $6.9/2\pi$ GHz. The frequency is given by $(\omega^{i}-\Omega)=(t^{i}-t_{\Omega})/\phi^{\prime\prime}$, where $t^{i}$ is the arrival time of the $i^{th}$ photon triggered by a sync signal from our Ti:sapphire laser and $\Omega=$193\text{\,}\mathrm{THz}$\$ is the central frequency of our 1550 nm photons. Since only the frequency difference $\Delta\omega^{i}$ is used in the MLE, we can set $t_{\Omega}=0$. The frequency of both output ports is then resolved as a linear function of the “time-of-flight” $\Delta\omega^{i}=\Delta t^{i}/\phi^{\prime\prime}$ through fibre with a dispersion factor $1/\phi^{\prime\prime}=0.0011(4)$ nm/ps.
SM3 Calibration and experimental parameters
@@
Table 1: Experimental parameters
Metric Gaussian $\sigma$ FWHM
Coherence length $\tau_{c}$ 1.24 2.92
Temporal bandwidth $\tau$ [ps] 0.98 2.31
Spectral bandwidth $\sigma/2\pi$ [GHz] 81 190
SM4 Estimation Protocol
@@
SM5 Estimator noise
Let the measured frequency shifts be $\{\Delta\omega_{i}\}_{i=1}^{N}$ with true density $P_{n}(\Delta\omega)$. We estimate $\Delta t$ by maximising the log likelihood $\ell(\Delta t)=\sum_{i=1}^{N}\log P(\Delta\omega_{i}\mid\Delta t)$, which is equivalent to using $P(\Delta t\mid\Delta\omega_{i})$ up to a normalising constant. Assume a small model mismatch between the true density $P_{n}$ and the model $P(\,\cdot\,\mid\Delta t)$. Write $\Delta\omega_{i}=\tilde{\Delta\omega}_{i}+\varepsilon_{i}$, where $\tilde{\Delta\omega}_{i}\sim P(\,\cdot\,\mid\Delta t)$ and $\varepsilon_{i}$ captures the difference between the two distributions with $\mathbb{E}[\varepsilon_{i}]=0$ and $\mathbb{E}[\varepsilon_{i}^{2}]=\sigma_{\varepsilon}^{2}$. A second-order Taylor expansion of each log likelihood term about $\tilde{\Delta\omega}_{i}$ gives,
$\log P(\Delta\omega_{i}\mid\Delta t)=\log P(\tilde{\Delta\omega}_{i}\mid\Delta t)+\partial_{\Delta\omega}\log P(\tilde{\Delta\omega}_{i}\mid\Delta t)\,\varepsilon_{i}+\tfrac{1}{2}\,\partial_{\Delta\omega}^{2}\log P(\tilde{\Delta\omega}_{i}\mid\Delta t)\,\varepsilon_{i}^{2}+o(\varepsilon_{i}^{2}).$ (SM2)
Taking expectations over the mismatch, the first order term vanishes since $\mathbb{E}[\varepsilon_{i}]=0$. The second order term contributes a bias proportional to $\mathbb{E}[\varepsilon_{i}^{2}]$, hence to $\sigma_{\varepsilon}^{2}$. This contribution increases as $P_{n}$ and $P(\,\cdot\,\mid\Delta t)$ diverge. In the regime $|\Delta\omega\,\Delta t|\ll 1$, the two distributions are practically equivalent, so $\sigma_{\varepsilon}^{2}$ is negligible and the induced bias can be ignored to an excellent approximation.
SM6 Spectral indistinguishability between one and two sources
@@
SM7 Stability of the setup over time
Collecting many samples over a long period of time can cause bias in the estimates due to experimental drift. One way to mitigate this is to analyse the Allan variance of the experimentally obtained samples for the bunching and coincidence outcomes, $\alpha(X)$, where X=1,1X=1,-1 and the spectral differences. In Fig. SM4 we have plotted the Allan variance for the BoC and spectral differences, both the calculated using the following form,
$\sigma_{y}^{2}(\tau)=\frac{1}{2\,(M-1)}\sum_{j=1}^{M-1}\bigl(\tilde{y}_{j+1}-\tilde{y}_{j}\bigr)^{2}.$ (SM3)
@@
SM8 Maximum resolution estimation
In this section, we analyse the resolution of the resolved optical delay estimation using our best prepared state. We use low pump power (100 mW) and achieve HOM visibility of 0.98, which improves the estimate precision. Fig. SM3 displays the cumulative estimate for the delays prepared at 6.567 ps and 6.570 ps.
[…]
SM9 Derivation of Probability function in terms of frequency difference
@@
$P_{\eta}(\omega,\omega^{\prime},X)=\gamma^{2}\xi(\omega)^{2}\xi(\omega^{\prime})^{2}\frac{1+\alpha(X)\eta^{2}\cos\left((\omega-\omega^{\prime})\Delta t\right)}{2},$ (SM4)
@@
$\xi(\omega)^{2}=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left[-\frac{\left(\omega-\Omega_{0}\right)^{2}}{2\sigma^{2}}\right].$ (SM5)
@@
$P_{\eta}(W,\Delta\omega,X)=\gamma^{2}\xi^{2}\left(W+\frac{\Delta\omega}{2}\right)\xi^{2}\left(W-\frac{\Delta\omega}{2}\right)\frac{1+\alpha(X)\eta^{2}\cos\left(\Delta\omega\Delta t\right)}{2}.$ (SM6)
@@
$\begin{split}\xi^{2}\left(W+\frac{\Delta\omega}{2}\right)\xi^{2}\left(W-\frac{\Delta\omega}{2}\right)&=\frac{1}{2\pi\sigma^{2}}\exp\left[-\frac{(W-\Omega_{0}+\Delta\omega/2)^{2}}{2\sigma^{2}}\right]\exp\left[-\frac{(W-\Omega_{0}-\Delta\omega/2)^{2}}{2\sigma^{2}}\right]\\ &=\frac{1}{2\pi\sigma^{2}}\exp\left[-\frac{(W-\Omega_{0})^{2}}{\sigma^{2}}\right]\exp\left[-\frac{\Delta\omega^{2}}{4\sigma^{2}}\right],\end{split}$ (SM7)